OpenStudy (anonymous):

on the soybean, cormeal percentage question how many pounds of cornmeal and how many pounds of soy beans gave a 12% mixture-exactly

OpenStudy (saifoo.khan):

incomplete question!

OpenStudy (anonymous):

could have sworn i just solved this one. maybe i should make a template, because it seems very popular

OpenStudy (anonymous):

you never mentioned exactly how many pounds of soybeans and cornmeal makes 12% in a 320 pound mixture.

OpenStudy (anonymous):

the web site you graciously gave me has far less a clue what the problem means than you and I do. Thanks anyway.

OpenStudy (saifoo.khan):

was my answer incorrect? :O

OpenStudy (anonymous):

I'll just guess it's about a 120 pounds of cornmeal and 200 pounds of soybeans and I'll probably be right.

OpenStudy (saifoo.khan):

was my answer incorrect?

OpenStudy (anonymous):

what was your answer ? I missed it.

OpenStudy (saifoo.khan):

Wow, great. it was 160 for both. hehe

OpenStudy (anonymous):

160 for both doesn't seem likely based on the amounts and pounds originally given.

OpenStudy (anonymous):

how did you come up with 160 for both?

OpenStudy (saifoo.khan):

Let s represent soybean meal and c cornmeal. s + c = 320 0.16s + 0.16c = 0.16 x 320 0.16s + 0.16c = 51.2 --------- 1 0.16s + 0.08c = 0.12 x 320 0.16s + 0.08c = 38.4 ---------- 2 Subtracting B from A: 0.08c = 12.8 c = 12.8/0.08 = 160 Hence s = 320 - 160 = 160 160 pounds of each should be mixed together in order to get a 320 pound mixture that is 12% protein.

OpenStudy (anonymous):

I know that if you have soybean that is 16% protien and cornmeal that is 8% you can not have an equal amount of both making 12% protien in a 120 pound batch.

OpenStudy (anonymous):

Maybe you didn't understand my question. My question was Soybean is 16% protien, cornmeal is 8% protien.How many pounds of each should be mixed together in order to get a 320 pound mixture that is 12% protien? This is not 50/50 answer